Music cd's are on sale. The original price was $ 20.00. The sale price is $ 25.00. What is the percent saving?
Answer by Fombitz(32388)

I think you mean the original price is $25, sale price is $20.
%=(New-Old)/Old=(20-25)/25=5/25=-20%
20% reduction in price.